# Data-driven model validation for neutrino-nucleus cross section measurements

### v1

@@ III.2 Developing a Model Validation Procedure
Given a flux prediction with its associated uncertainties, the correlations that yield $\mu^{E_{\mathrm{had}}^{\mathrm{rec}},\text{const.}}$ will be dictated by the modeling of $E_{\mathrm{had}}^{\mathrm{missing}}$. Thus, if $E_{\ell}^{\mathrm{rec}}$ and $E_{\mathrm{had}}^{\mathrm{rec}}$ are measured, and a constraint from $E_{\ell}^{\mathrm{rec}}$ is applied to $E_{\mathrm{had}}^{\mathrm{rec}}$ to reduce uncertainties on the flux and overall $E_{\nu}$ prediction, a precise description of $E_{\mathrm{had}}^{\mathrm{missing}}$ becomes necessary for Eq. 12 to be satisfied. It is this narrowed parameter space for the correlated $E_{\ell}$, $E_{\mathrm{had}}^{\mathrm{vis}}$, and $E_{\mathrm{had}}^{\mathrm{missing}}$ predictions that provide sensitivity to mis-modeling in the missing energy through conservation of energy.
@@ IV.2.1 Proton Energy Scaling Fake Data Studies
Table 1: Results of the model validation and cross-section extraction for each fake data set that scales the reconstructed proton energy. All sources of uncertainty are included at data statistics equivalent to 6.4×$\times 10^{20}$ protons on target (POT) of exposure. The significance values in the “Bias in Extraction” columns indicate the level of closure the extracted fake data cross-sections have with the underlying truth. The significance values in the “Sensitivity of the Model Validation to Discrepancies” columns indicate the level at which the fake data and the MC used for the cross section extraction disagree. The different columns correspond to different validation tests.
@@
Encountering a situation in real data analogous to the multivariate event reweighter, which would pass the model validation but has an entirely different mapping between true and reconstructed neutrino energy, cannot be completely ruled out. However, such a scenario is unlikely as lepton kinematic distributions are well constrained by electron scattering data, particularly for QE events. The BDT reweighting we explore here shows noticeable tension with the QE prediction from the unreweighted model and performs worse when compared to the QE-like $\nu_{\mu}$1p dataset from [47, 48]. Given the extremity of the example, this DUNE FDS is not inconsistent with the notion that one is able to use data-driven validation to detect mis-modeling of the mapping between true and reconstructed neutrino energy to enable the extraction of energy dependent cross sections.
@@
Table 2: Results of the model validation and cross section extraction for the NuWro FDSs. Two treatments of systematic uncertainties are considered: one with full systematic uncertainties (“All Uncertainties”) and the other with only cross section induced systematic uncertainties (“XS Syst+Stat Uncertainties”). Statistical uncertainties are always included. The Bias in Extraction columns indicate the significance of the tension between the extracted result and NuWro prediction. The entries in the Model Validation columns indicate the significance at which the fake data and the MC used for the cross section extraction disagree in the test that shows the most tension. These significance levels were computed from the $p$-values obtained in each test or cross section extraction.
@@ IV.2.2 Alternative Event Generator Fake Data Studies
Figure 13: The (a) NuWro and (b) MicroBooNE MC $\cos\theta_{\mu,p}^{\mathrm{rec}}$ predictions broken down by interaction mode.
@@
Figure 14: Model validation test comparing the NuWro fake data sample to the nominal MC prediction for the reconstructed muon-proton opening angle distribution for FC events. The red (blue) lines and bands show the prediction without (with) the constraint from the reconstructed muon kinematics. The uncertainties of the prediction only include the cross section and statistical terms and are shown in the bands. The statistical uncertainties on the data are shown on the data points.
[…]
@@
To further explore the behavior of the conditional constraint from the muon kinematics and its potential utility beyond its role in model validation for cross section extraction, we investigate its impact on the distribution of the reconstructed muon-proton opening angle, $\theta_{\mu,p}^{\mathrm{rec}}$. This kinematic variable is sensitive to the breakdown of different interaction modes with QE interactions dominating at negative $\cos\theta_{\mu,p}^{\mathrm{rec}}$ and MEC and RES events more prominent at positive $\cos\theta_{\mu,p}^{\mathrm{rec}}$. Figure 13, which displays the NuWro and MicroBooNE MC $\cos\theta_{\mu,p}^{\mathrm{rec}}$ predictions broken down by interaction mode, illustrates this separation. The significantly different MEC predictions from the two generators is also apparent here. This difference is still present when the NuWro fake data $\cos\theta_{\mu,p}$ distribution is examined after a constraint from the muon kinematics, as can be seen in Fig. 14. Agreement in the more QE-rich region ($\cos\theta_{\mu,p}^{\mathrm{rec}}<0$) is observed after constraint whereas the MEC-rich region ($\cos\theta_{\mu,p}^{\mathrm{rec}}>0.5$) shows noticeable discrepancies. Given the similar RES predictions between the two models, this suggests that the MEC prediction from NuWro is outside the uncertainties of the MicroBooNE model. This study demonstrates how the model validation and the conditional constraint can be used to probe the modeling of the relative contribution from different interaction modes.
@@ V Summary
The heavy reliance on event generators makes model validation important when extracting neutrino-nucleus interaction cross sections. To this end, we have presented a set of data-driven validation procedures and demonstrated that they can be powerful for detecting deficiencies in the models used for cross-section extraction. This validation utilizes techniques based on goodness-of-fit tests and the conditional constraint procedure to determine if the overall model can describe the data within uncertainties in the phase space relevant for the unfolding. Since this approach is based upon real data, it is well-grounded to appropriately validate the unfolding model and to evaluate the need for additional uncertainties. If a carefully selected set of data-driven tests are passed by the model, this builds confidence that any bias introduced in the cross section extraction will be within the quoted uncertainties of the measurement. This applies to both measurements of visible kinematic variables, such as the outgoing muon kinematics, and to derived quantities, such as the energy transferred to the nucleus. We demonstrate the efficacy of these data-driven methods with fake data studies aimed at comparing the sensitivity of the validation to the amount of bias introduced in the cross section extraction, in which we find that the validation is able to detect mismodeling before it impacts the cross section extraction.

### v2

@@ III.2 Developing a Model Validation Procedure
Given a flux prediction with its associated uncertainties, the correlations that yield $\mu^{E_{\mathrm{had}}^{\mathrm{rec}},\text{const.}}$ will be dictated by the modeling of $E_{\mathrm{had}}^{\mathrm{missing}}$. Thus, if $E_{\ell}^{\mathrm{rec}}$ and $E_{\mathrm{had}}^{\mathrm{rec}}$ are measured, and a constraint from $E_{\ell}^{\mathrm{rec}}$ is applied to $E_{\mathrm{had}}^{\mathrm{rec}}$ to reduce uncertainties on the overall $E_{\nu}$ prediction through the high correlation in the flux prediction shared between the channels, a precise description of $E_{\mathrm{had}}^{\mathrm{missing}}$ becomes necessary for Eq. 12 to be satisfied. It is this narrowed parameter space for the correlated $E_{\ell}$, $E_{\mathrm{had}}^{\mathrm{vis}}$, and $E_{\mathrm{had}}^{\mathrm{missing}}$ predictions that provide sensitivity to mis-modeling in the missing energy through conservation of energy. Nevertheless, it should be noted that Eq. 12 still allows for a conspiracy of flux mis-modeling (aka mis-modeling of $E_{\nu}$) and mis-modeling of $E_{\mathrm{had}}^{\mathrm{missing}}$ to cancel out and pass the test described by Eq. 14. However, the use of the constraint brings the examination of $E_{\mathrm{had}}^{\mathrm{missing}}$ through the correlations between $E_{\ell}^{\mathrm{rec}}$ and $E_{\mathrm{had}}^{\mathrm{rec}}$ onto more equal footing with the examination of a directly observable variable, like $E_{\ell}^{\mathrm{rec}}$, in an unknown neutrino flux spectrum. As in the case of $E_{\mathrm{had}}^{\mathrm{missing}}$, an $E_{\mu}$ prediction is always going to depend on the assumed $E_{\nu}$ and it is challenging to fully disentangle the two.
@@ IV.2.1 Proton Energy Scaling Fake Data Studies
Table 1: Results of the model validation and cross-section extraction for each fake data set that scales the reconstructed proton energy. All sources of uncertainty are included at data statistics equivalent to 6.4×$\times 10^{20}$ protons on target (POT) of exposure. The significance values in the “Bias in Extraction” columns indicate the level of closure the extracted fake data cross-sections have with the underlying truth. The significance values in the “Sensitivity of the Model Validation to Discrepancies” columns indicate the level at which the fake data and the MC used for the cross section extraction disagree. The different columns correspond to different validation tests. All significance levels were obtained by converting the $\chi^{2}$ test statistic to a $p$-value assuming a $\chi^{2}$ distribution with degrees of freedom equal to the number of bins. These $p$-values are then interpreted as significance levels corresponding to multiples of the standard deviation $\sigma$ of a normal distribution.
@@
A comparison between the weights obtained in this study and those obtained in the DUNE FDS can be seen in Fig. 8. Similar behavior is observed with QE events at high $Q^{2}$ scaled up by as much as a factor of 2 and QE events at low $Q^{2}$ scaled down by as much as a factor of 0.6. As such, when the true $Q^{2}$ distribution for QE events is compared to the nominal model, a significant discrepancy is seen. This is demonstrated in Fig. 9, where the $\chi^{2}/\mathrm{ndf}$ calculated between the reweighted distribution and the nominal model and its associated cross section uncertainties is $135.04/25$, indicating significant discrepancy. The basic lepton kinematics distributions for QE events are agnostic to final state interactions in the limit of complete factorization of the initial and final interactions, which is a well realized approximation employed by state-of-the-art neutrino event generators [18, 51]. As such, being supported by many neutrino and electron scattering experiments, these distributions are one of the better understood portions of cross section modeling, making such a large discrepancy with the nominal model hard to justify.
@@
Encountering a situation in real data analogous to the multivariate event reweighter, which would pass the model validation but has an entirely different mapping between true and reconstructed neutrino energy, cannot be completely ruled out. However, such a scenario is unlikely as lepton kinematic distributions are well constrained by electron scattering data, particularly for QE events. The BDT reweighting we explore here shows noticeable tension with the QE prediction from the unreweighted model and performs worse when compared to the QE-like $\nu_{\mu}$1p dataset from [52, 53]. The DUNE study demonstrates that, through the use of PRISM, one can still obtain a robust measurement of oscillation parameters even in such a scenario with large deviations from reasonable models, which would be required if one wished to perform an oscillation analysis with minimal assumptions on the cross section model. However, techniques such as PRISM are not applicable to many experiments, such as MicroBooNE, and other strategies like data-driven model validation are required. Likewise, we caution that the general conclusions about the model validation drawn from the FDSs present in this work may not extrapolate as well onto scenarios far outside our current understanding of neutrino-nucleus scattering. The same holds true if more “tradition” FDSs are used to validate a model for a cross section measurement; one should be cautious extrapolating the results of successful closure on alternative models well beyond the bounds of reasonable models if these alternative models do not also extend well beyond these bounds. Indeed, this is directly related to the three issues raised earlier on “tradition” FDSs. There is no guarantee that a model spread will cover nature (issue 1), but including too many extreme scenarios may result in overestimated uncertainties (issue 2), and it is not clear when one should stop examining closure on these alternative models, particularly when extreme deviations are included (issue 3). Given the extremity of the example, this DUNE FDS is not inconsistent with the notion that one is able to use data-driven validation to detect mis-modeling of the mapping between true and reconstructed neutrino energy to enable the extraction of energy dependent cross sections.
@@
Table 2: Results of the model validation and cross section extraction for the NuWro FDSs. Two treatments of systematic uncertainties are considered: one with full systematic uncertainties (“All Uncertainties”) and the other with only cross section induced systematic uncertainties (“XS Syst+Stat Uncertainties”). Statistical uncertainties are always included. The Bias in Extraction columns indicate the significance of the tension between the extracted result and NuWro prediction. The entries in the Model Validation columns indicate the significance at which the fake data and the MC used for the cross section extraction disagree in the test that shows the most tension. The significance levels were obtained using $p$-values calculated from the $\chi^{2}$ test statistic interpreted assuming a $\chi^{2}$ distribution with degrees of freedom equal to the number of bins.
@@ IV.2.2 Alternative Event Generator Fake Data Studies
Figure 13: The (a) NuWro and (b) MicroBooNE MC $\cos\theta_{\mu,p}^{\mathrm{rec}}$ predictions for events passing the semi-inclusive $\nu_{m}u$CCNp selection broken down by interaction mode.
Figure 14: Model validation test comparing the NuWro fake data sample to the nominal MC prediction for the reconstructed muon-proton opening angle distribution formed from events passing the semi-inclusive $\nu_{\mu}$p selection. The red (blue) lines and bands show the prediction without (with) the constraint from muon kinematics for the distribution formed from events passing the $\nu_{\mu}$p selection and the distribution formed from events passing the $\nu_{\mu}$0p selection. The uncertainties of the prediction include all systematic and statistical terms. These are shown in the bands on the prediction. The statistical uncertainties on the fake data are shown on the data points.
NuWro Fake Data MicroBooNE MC MicroBooNE MC with constraint
QE 0.449 0.366$\pm$0.058 0.456$\pm$0.025
MEC 0.116 0.218$\pm$0.104 0.145$\pm$0.032
RES+DIS 0.375 0.367$\pm$0.093 0.352$\pm$0.035
Table 3: Fractional contribution of $\nu_{\mu}$ events from different interaction modes to the total $\nu_{\mu}$ Np selection. The remaining $\sim$5of events are mostly background non-$\nu_{\mu}$ events that pass the selection.
@@
IV.2.3 Further exploration of the conditional constraint
@@
In what follows, we explore the potential utility of conditional constraints beyond their role in model validation for cross section extraction. To do so, we further investigate the behavior of the constraint from the muon kinematics by examining its impact on the distribution of the reconstructed muon-proton opening angle, $\theta_{\mu,p}^{\mathrm{rec}}$, for a semi-inclusive selection of $\nu_{\mu}$ events with protons ($\nu_{\mu}$ Np). This selection is identical to the $\nu_{\mu}$ Np selection described in Ref. [12] and consists of all events passing the inclusive $\nu_{\mu}$ selection that also have a primary reconstructed proton with greater than 35 MeV of kinetic energy. The $\theta_{\mu,p}^{\mathrm{rec}}$ distribution is particularity interesting because it is sensitive to the breakdown of different interaction modes, with QE interactions dominating at negative $\cos\theta_{\mu,p}^{\mathrm{rec}}$ and MEC and RES+DIS events more prominent at positive $\cos\theta_{\mu,p}^{\mathrm{rec}}$. Figure 13, which displays the NuWro and MicroBooNE MC $\cos\theta_{\mu,p}^{\mathrm{rec}}$ predictions broken down by interaction mode, illustrates this separation.
Figure 13 illustrates that there is a noticeable difference in the strength of the MEC predictions from NuWro and the MicroBooNE MC. In addition, the QE prediction from the MC is somewhat larger than NuWro, while the RES+DIS contributions are similar. This is quantified in Table. 3, which contains the fraction contribute to the total selection for $\nu_{\mu}$ events from each interaction mode. These differences are also apparent when the NuWro fake data $\cos\theta_{\mu,p}$ distribution is compared to the MicroBooNE MC prediction with all its associated uncertainties. This is shown in Fig. 14. Even with the full uncertainty band, the more MEC and RES+DIS rich region ($\cos\theta_{\mu,p}^{\mathrm{rec}}>0$) is overpredicted to the point that the NuWro fake data is not completely covered by the uncertainties on prediction. The fake data is also overpredicted in the QE rich region ($\cos\theta_{\mu,p}^{\mathrm{rec}}<0$), but to a lesser extent and still falls within the uncertainties of the MC.
To further examine the data to MC differences in this observable, we expand the typical procedure of applying the constraint to the total prediction by also appling it to the prediction for each individual interaction mode. The constraint utilizes the muon kinematics distributions of the FC&PC $\nu_{\mu}$p selection and the FC&PC $\nu_{\mu}$0p selection, which consists of the subset of events in the inclusive $\nu_{\mu}$ selection without a final state proton or with a final state proton below the 35 MeV threshold. The latter selection is likewise equivalent to the $\nu_{\mu}$0p selection described in Ref. [12]. This application of the constraint to individual interaction modes rather than the prediction as a whole is only done in this study for the purposes of interpreting the action of the constraint, this procedure is not used in any of the model validation described throughout the rest of this work. As expected based on the difference in size of the MEC contribution between the two generators, the constraint produces a noticeable decrease in the MicroBooNE MC MEC prediction. A decrease in the RES+DIS prediction is also observed. Nevertheless, the fake data still falls at the lower edge of the resulting total posterior prediction’s uncertainties for $\cos\theta_{\mu,p}^{\mathrm{rec}}>0.5$ where the contribution from QE interaction modes is small and MEC and RES+DIS interaction modes dictate the strength of the prediction. For $\cos\theta_{\mu,p}^{\mathrm{rec}}<0.5$, a small increase in the QE prediction offsets the decrease in the other interaction modes. This causes the total prediction to decrease slightly such that it agrees with the fake data within uncertainties across the QE dominated region of phase space.
Quantitatively, the posterior QE prediction corresponds to 0.456$\pm$0.025 of the selection. This is in good agreement with the NuWro fake data, for which 0.449 of the selection is from QE $\nu_{\mu}$ interactions. This indicates that the constraint has utilized the information provided by the muon kinematics, with the only hadronic information provided by splitting the distribution based on the presence of a $>35$ MeV kinetic energy proton, to increase the QE such that it falls in line with the NuWro fake data. Similarly, the prediction for the fraction of MEC events decreases from 0.218$\pm$0.104 to 0.145$\pm$0.032 and it also agrees with the fake data for which the fraction of MEC events is 0.116. However, despite the initially similar RES+DIS agreement, the constraint reduces this component such that the posterior prediction of 0.352$\pm$0.035 is further from the NuWro value of 0.375, but is nevertheless still within uncertainties. These metrics are summarized in Table. 3.
This study demonstrates how the model validation and the conditional constraint can be used to probe the modeling of different interaction modes and their composition. The constraint greatly reduces the uncertainties on the rates ($>50\%$) for each mode and tends to brings their CVs closer to the truth from the fake data. More generally, this study demonstrates how these methods allow one to further examine where the data lies within the allowed model space with significantly more sensitivity. The versatility of these methods means they can be utilized to provide additional insight about the models in such a way to be complementary to the actual cross section extraction. This opens up additional avenues of investigation for future analyses. For example, one could use the conditional constraint to fit various free parameters within a given model more directly. This can be easily achieved by denoting the free parameter as the constrained variable and using the distribution of a given observable as the constraining channel. From then, the rest of the procedure follows from Eqs. 10 and 11, which updates the free parameter to its posterior prediction based on an assumed prior and the given model. However, we note that interpreting these results in this way is by construction, model dependent and care must be taken to do so across different models, given that there is likely a complex interplay between implementation of these free parameters and other parameters of that model. Nevertheless, this can be quite enlightening and also complementary to the extracted cross section which, though potentially not providing as direct of insight, describes a well-defined physics quantity and will be interpretable in the context of any other model prediction.
@@ V Summary
The heavy reliance on event generators makes model validation important when extracting neutrino-nucleus interaction cross sections. To this end, we have presented a set of data-driven validation procedures and demonstrated that they can be powerful for detecting deficiencies in the models used for cross-section extraction. This validation utilizes techniques based on goodness-of-fit tests and the conditional constraint procedure to determine if the overall model can describe the data within uncertainties in the phase space relevant for the unfolding. Since this approach is based upon real data, it is well-grounded to appropriately validate the unfolding model and to evaluate the need for additional uncertainties. If a carefully selected set of data-driven tests are passed by the model, this builds confidence that any bias introduced in the cross section extraction will be within the quoted uncertainties of the measurement. This applies to both measurements of visible kinematic variables, such as the outgoing muon kinematics, and to derived quantities, such as the energy transferred to the nucleus. We demonstrate the efficacy of these data-driven methods with fake data studies aimed at comparing the sensitivity of the validation to the amount of bias introduced in the cross section extraction, in which we find that the validation is able to detect mis-modeling before it impacts the cross section extraction. These data-driven methods based upon the conditional constraint are highly versatility and can be utilized to provide additional insight that complements the actual cross section extraction.